OpenStudy (anonymous):

a circle has the equation 4x^2-32+4y^2=16 find the center (h,k) and the radius

OpenStudy (anonymous):

Write it as \[ 4x^2-32+4y^2=16\\ 4x^2+4y^2=48\\ x^2 + y^2=21 \] You should be able to find the answers now. Can you?

OpenStudy (mertsj):

@eliassaab You might want to check your division

OpenStudy (anonymous):

yeah it should be x^2+y^2=3

OpenStudy (anonymous):

i mean 12

OpenStudy (anonymous):

Oh \[ x^2 + y^2=12 \]

OpenStudy (anonymous):